# MAPGD: Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization

### v2

@@ 3.3. Hypersphere-Constrained Gradient Clustering
Details of Hypersphere-Constrained Gradient Clustering: Intra-cluster Compactness and Inter-cluster Repulsion, Original Boundary, Angular Margin
Optimization Objective.
Analogous to hyperspherical contrastive learning, the clustering objective combines intra-cluster compactness and inter-cluster repulsion:
$\mathcal{L}_{\text{HCGC}}=-\log\frac{\exp(\cos(\Delta(\hat{v}_{i},\hat{v}_{j}))/\tau)}{\sum_{k\in\mathcal{N}(i)}\exp(\cos(\Delta(\hat{v}_{i},\hat{v}_{k}))/\tau)},$ (7)
where $(\hat{v}_{i},\hat{v}_{j})$ is a positive pair from the same cluster, $\mathcal{N}(i)$ denotes negatives from other clusters, and $\tau$ is a temperature factor. By optimizing $\mathcal{L}_{\text{HCGC}}$, we obtain semantically well-aligned clusters of gradients, ensuring coherent downstream fusion.
After clustering, gradients within each semantically coherent group are fused into a single representative gradient. This fusion is typically performed by an LLM, guided by the original suggestions. To maintain diversity and avoid redundancy among the fused gradients, a semantic diversity filter is applied, discarding suggestions that are too similar to already selected ones, controlled by a diversity threshold $\theta_{diversity}$.
Figure 4. Illustration of agent-generated pseudo-gradients in a sentiment classification task. Blue gradients form a coherent cluster under the Instruction Specialist, while the red gradient represents a conflicting update. Green gradients from the Example Curator constitute another semantically compact cluster. This motivates the need for hypersphere-constrained clustering to enforce intra-cluster compactness and inter-cluster separation.
@@
Illustration of agent-generated pseudo-gradients in a sentiment classification task. Blue gradients form a coherent cluster under the Instruction Specialist, while the red gradient represents a conflicting update. Green gradients from the Example Curator constitute another semantically compact cluster. This motivates the need for hypersphere-constrained clustering to enforce intra-cluster compactness and inter-cluster separation.
3.4. Channel-Adaptive Agent Weighting
While HCGC ensures structural consistency across gradients, gradients within the same cluster may still differ in reliability due to heterogeneous agent expertise. To address this, we propose Channel-Adaptive Agent Weighting (CAAW), which dynamically calibrates agent contributions according to their historical effectiveness, akin to channel-wise recalibration in neural representations Hu et al. (2018).
Adaptive Weight Assignment.
Let $s_{k}^{(t)}$ denote the validation performance gain associated with gradient $g_{k}^{(t)}$. The adaptive weight of each gradient is computed via a softmax normalization:
where $C$ is the current cluster, and $\lambda$ controls the sharpness of weighting. This formulation emphasizes historically effective agents while down-weighting less reliable ones.
Gradient Fusion.
The final fused gradient is expressed as:
$g_{\text{fused}}^{(t)}=\Psi\left(\sum_{k\in C}w_{k}^{(t)}g_{k}^{(t)}\right),$ (9)
where $\Psi(\cdot)$ denotes an LLM-driven synthesis function that integrates multiple textual directions into a coherent refinement.
Robustness and Stability.
By recalibrating agent contributions at the channel level, CAAW suppresses redundant or conflicting updates while amplifying stable and discriminative directions. Theoretically, this reduces variance in the optimization trajectory and prevents over-amplification of noisy gradients, thereby ensuring convergence to semantically robust prompt candidates.
3.5. Candidate Generation and Selection
Fused gradients generate successor prompts $\{p^{\prime}_{1},\dots,p^{\prime}_{n}\}$. MAPGD applies lightweight filtering and optional bandit-based evaluation Audibert and Bubeck (2010) to control computational overhead. However, unlike prior work that heavily relies on bandit selection, our framework emphasizes geometry-aware clustering (HCGC) and adaptive agent weighting (CAAW) as the primary mechanisms ensuring robustness and interpretability.
3.6. Theoretical Convergence (Informal)
We provide a convergence analysis of MAPGD in Appendix C. Here we summarize the key assumptions: (i) bounded variance of agent pseudo-gradients, (ii) Lipschitz continuity of the loss function, and (iii) unbiased sampling from bandit-based candidate selection. Under these conditions, MAPGD inherits the convergence properties of stochastic gradient descent and achieves almost sure convergence to a local optimum at a sublinear rate of $\mathcal{O}(1/\sqrt{T})$. Intuitively, HCGC reduces variance by enforcing geometric separation, CAAW emphasizes historically reliable agents, and bandit selection maintains unbiased exploration, together ensuring stable optimization dynamics.
4. Experiments
We design comprehensive experiments to evaluate the effectiveness, efficiency, and robustness of the proposed MAPGD framework. To ensure comparability, we reproduce the experimental setup of ProTeGi under identical conditions and extend the evaluation with ablation studies and token consumption analysis. In addition, we include a real-world case study on prompt optimization for text generation (see Appendix D), which highlights the applicability of MAPGD beyond controlled benchmarks.
4.1. Datasets
@@
4.2. Experimental Setup
For each dataset, we randomly sample 50 examples for development and 150 for testing. All reported results are averaged over three independent runs to mitigate variance. The evaluation metric is binary F1 score on the test set, computed by max-pooling over the final beam of candidate prompts.
Unless otherwise stated, all experiments are conducted with the January 2023 version of gpt-3.5-turbo through the Azure OpenAI API. The temperature is set to 0.0 for classification tasks to ensure deterministic predictions.
To ensure a fair comparison, all methods—including baselines and MAPGD—use the same initial prompts, training data, and random seeds. The iteration budget is fixed at $T=10$ across all experiments. We adopt default parameter values without additional hyperparameter search, to ensure that performance differences are attributable to algorithmic design rather than parameter tuning.
For our MAPGD implementation, we instantiate the framework with specific hyperparameters introduced in Section 3.3. We use the all-MiniLM-L6-v2 Sentence Transformer as the gradient encoder $\phi$. The conflict threshold $\theta_{conflict}$ is set to 0.3, and we use a cluster similarity threshold of 0.7 for grouping gradients. The diversity threshold $\theta_{diversity}$ is set to 0.7 to ensure varied candidate prompts. Bandit-based selection adopts a UCB strategy with a budget of 80 evaluations.We set the CAAW weighting parameter $\lambda$ to 1 (see Appendix A for a sensitivity analysis on different values).
Accordingly, all experiments reported in Section 4.4 were conducted with these modules enabled, unless explicitly ablated.
4.3. Baselines
@@
4.4. Experimental Results
Figure 5 summarizes the overall performance across the four benchmark classification datasets. Our results demonstrate that MAPGD consistently achieves the best performance under varying evaluation budgets, outperforming all baselines by a notable margin.
On the Jailbreak dataset, MAPGD attains an average F1 of 0.86, surpassing ProTeGi (0.82) and MC (0.76), while substantially improving over RL (0.67) and AutoGPT (0.62). The improvement margin remains stable across different budgets, reaching up to +5.5% over ProTeGi and +10.9% over MC at 30 evaluations.
For Ethos, MAPGD pushes performance from the strong baseline of 0.93 ($p_{0}$) to 0.98, showing a consistent +1.5% advantage over ProTeGi and +4.0% over MC. These improvements highlight MAPGD’s ability to leverage gradient fusion even in relatively saturated tasks.
On the more challenging LIAR dataset, MAPGD achieves 0.71, representing the largest relative gains: +8.0% over ProTeGi, +11.0% over MC, and a substantial +17.0% over the initial prompt. This suggests that multi-agent collaboration and semantic clustering are particularly beneficial in noisy, fact-checking tasks.
For Sarcasm, MAPGD steadily improves from 0.84 ($p_{0}$) to 0.91, outperforming ProTeGi (0.87) and MC (0.85). Interestingly, AutoGPT sometimes reduces performance below the initial prompt, echoing earlier findings that unguided agent feedback may destabilize optimization.
Beyond classification, we further evaluate MAPGD on arithmetic reasoning benchmarks, which require detailed multi-step inference. Table 1 compares MAPGD against three strong baselines—InsZero, Instinct, and PromptWizard—on GSM8k, AQUARAT, and SVAMP. MAPGD achieves consistent improvements across all datasets, with particularly notable gains on GSM8k (93.5%), surpassing the previous best by +3.5 points. On AQUARAT and SVAMP, MAPGD reaches 60.3% and 84.1%, respectively, representing stable improvements over the strongest baselines. These results confirm that MAPGD not only excels in classification but also transfers effectively to challenging reasoning tasks.
Overall, MAPGD improves over ProTeGi and MC by 3–7% on average for classification, while maintaining strong gains on arithmetic reasoning datasets. These results validate our hypothesis that multi-agent specialization and hypersphere-constrained clustering provide more robust and adaptive optimization compared to single-agent or rule-based approaches.
Table 1. Arithmetic reasoning accuracy (%) on GSM8k, AQUARAT, and SVAMP. Best results in bold.
Method GSM8k AQUARAT SVAMP
InsZero 74.2 54.3 79.5
Instinct 74.5 54.7 81.0
PromptWizard 90.0 58.2 82.3
MAPGD 93.5 60.3 84.1
Figure 5. Test performance (F1 score) vs. API query budget per prompt candidate across four benchmark tasks.
[…]
Test performance (F1 score) vs. API query budget per prompt candidate across four benchmark tasks.
4.5. Ablation Study
To precisely attribute the performance gains of MAPGD, we perform a comprehensive ablation study (see Table 2 for the detailed configurations and results). Beyond standard ablation variants (e.g., No Gradient Fusion, Single-Agent Only, No Specialization), we design targeted experiments specifically to isolate the contributions of the two proposed modules HCGC and CAAW, as well as to evaluate resource efficiency and robustness.
Ablation configurations.
We evaluate the following variants:
  1. (1)
    MAPGD (full): our full method with HCGC and CAAW.
  2. (2)
    No Gradient Fusion: disable the fusion stage; agents act independently and their suggestions are applied without semantic clustering.
  3. (3)
    Single-Agent Only: only one specialized agent (instruction specialist) is used.
  4. (4)
    No Specialization: agent roles are randomized each iteration (no fixed expertise).
  5. (5)
    w/o HCGC: keep clustering but remove the hypersphere angular-margin constraint. This reduces to standard K-means clustering with cosine similarity, which acts as an off-the-shelf baseline for comparison.
  6. (6)
    w/o CAAW: replace Channel-Adaptive Agent Weighting with simple uniform averaging of agent gradients.
  7. (7)
    w/o HCGC+CAAW: use standard K-means clustering with cosine similarity (no angular margin) and replace adaptive weighting with uniform fusion.
Table 2. Ablation study: classification performance (F1 score, mean $\pm$ std over 3 runs). Best results in bold.
Method Jailbreak Ethos LIAR Sarcasm
MAPGD (full) 0.88 $\pm$ 0.01 0.98 $\pm$ 0.02 0.71 $\pm$ 0.01 0.91 $\pm$ 0.01
No Gradient Fusion 0.83 $\pm$ 0.02 0.92 $\pm$ 0.02 0.67 $\pm$ 0.02 0.86 $\pm$ 0.02
Single Agent Only 0.80 $\pm$ 0.03 0.87 $\pm$ 0.03 0.62 $\pm$ 0.02 0.82 $\pm$ 0.03
No Specialization 0.80 $\pm$ 0.02 0.90 $\pm$ 0.02 0.65 $\pm$ 0.02 0.84 $\pm$ 0.02
w/o HCGC 0.84 $\pm$ 0.02 0.94 $\pm$ 0.02 0.68 $\pm$ 0.02 0.87 $\pm$ 0.02
w/o CAAW 0.85 $\pm$ 0.02 0.95 $\pm$ 0.02 0.69 $\pm$ 0.02 0.88 $\pm$ 0.02
w/o HCGC + CAAW 0.82 $\pm$ 0.03 0.91 $\pm$ 0.03 0.66 $\pm$ 0.02 0.83 $\pm$ 0.02
4.6. Token Consumption Analysis
To further assess the efficiency of MAPGD, we conducted a dedicated token consumption experiment, following the same setup as in Section 4.2. In this experiment, we recorded the number of tokens used in every API call, where the token cost of a single call is defined as the sum of input tokens and output tokens. For each method, we accumulated token usage across all iterations and runs, and report the total tokens, average tokens per iteration, and total number of LLM calls. In addition, we report performance efficiency as the ratio of average F1 score to total token usage.
Table 3 presents the results, comparing MAPGD with the ProTeGi. MAPGD not only achieves higher F1 scores, but also requires fewer tokens and API calls overall. Specifically, MAPGD reduces the average token cost per iteration from roughly 25.6k to 23.6k, corresponding to a $\sim$8% saving in token usage. At the same time, MAPGD improves the performance-per-token efficiency by more than 10%. These findings indicate that multi-agent collaboration with HCGC and CAAW enables more effective use of query budgets, reducing redundancy while preserving accuracy.
Table 3. Token consumption comparison between ProTeGi and MAPGD (averaged over 3 runs). MAPGD achieves lower token usage and higher efficiency.
Metric ProTeGi MAPGD (full)
Total Tokens 256k 236k
Calls 962 643
Avg F1 0.83 0.87
F1 / Token ($\times 10^{-6}$) 3.24 3.69
4.7. Agent Number Sensitivity
To better justify the design choice of using four specialized agents, we conduct a sensitivity analysis by varying the number of agents $N\in\{2,4,6\}$. For $N=2$, we retain only the Instruction Specialist and Example Curator; for $N=6$, we extend the framework with two additional stylistic and format-focused agents. All other settings follow Section 4.2.
Figure 6 shows the average F1 score across the four datasets. We observe that performance improves substantially when moving from $N=2$ to $N=4$, confirming the benefit of multi-agent specialization. Increasing to $N=6$ yields only marginal gains while incurring higher token cost, suggesting diminishing returns. These results indicate that four agents provide a favorable balance between effectiveness and efficiency, validating our default configuration.
Overall, these findings suggest that an excessively small agent set cannot comprehensively capture different prompt dimensions, whereas an overly large set introduces redundancy and potential conflicts. A moderate number of agents, four in our case, provides the best trade-off.
Figure 6. Agent number sensitivity analysis. Increasing agents from 2 to 4 yields clear improvements, while further increasing to 6 brings diminishing returns.
[…]
Agent number sensitivity analysis. Increasing agents from 2 to 4 yields clear improvements, while further increasing to 6 brings diminishing returns.
4.8. Experimental Insights
Our experimental evaluation provides several key insights:
  • Effectiveness of multi-agent collaboration. Across all four benchmark tasks, MAPGD consistently outperforms single-agent ProTeGi and other baselines. The largest gains are observed on LIAR and Jailbreak, where conflicting gradient signals are common. This demonstrates that multi-agent specialization, combined with semantic clustering, can effectively resolve gradient conflicts and guide prompt optimization toward more robust improvements.
  • Contributions of HCGC and CAAW. The ablation study confirms that both proposed modules are indispensable. Removing HCGC reduces the ability to enforce semantic compactness and separation, leading to less stable optimization. Disabling CAAW degrades robustness by failing to downweight weaker agents. When both modules are removed, MAPGD collapses close to single-agent performance, highlighting their complementary roles.
  • Efficiency in token usage. Despite employing multiple agents, MAPGD achieves lower overall token consumption and fewer API calls than ProTeGi. This efficiency stems from clustering redundant gradients and adaptively weighting agents, which reduce unnecessary evaluations. As a result, MAPGD achieves over 10% higher performance-per-token efficiency, indicating that it not only improves accuracy but also makes more economical use of query budgets.
  • Robustness across tasks. The improvements are consistent across datasets of varying size, domain, and language, suggesting that MAPGD generalizes well to heterogeneous NLP tasks. Furthermore, the case study in the Appendix illustrates its applicability to open-ended generation, extending beyond controlled classification benchmarks.
  • Transferability to reasoning tasks. MAPGD also demonstrates strong performance on arithmetic reasoning datasets, surpassing competitive baselines such as PromptWizard. Its consistent improvements on GSM8k, AQUARAT, and SVAMP highlight that the proposed gradient-inspired multi-agent optimization is not limited to classification, but can effectively enhance chain-of-thought style reasoning as well.
In summary, these insights show that MAPGD achieves a favorable balance between performance and efficiency, while its modular design (HCGC + CAAW) provides both effectiveness and robustness in diverse prompt optimization scenarios.
5. Discussion
MAPGD advances prompt optimization by combining gradient-inspired reasoning with multi-agent collaboration. Its two key mechanisms—Hypersphere-Constrained Gradient Clustering and Channel-Adaptive Agent Weighting enable more structured, interpretable, and efficient optimization.
Modularity. By separating gradient generation, clustering, weighting, and candidate selection, MAPGD clarifies which components drive performance, supporting transparent analysis and principled extensions.
Conflict resolution. HCGC enforces angular margin constraints on the hypersphere, promoting intra-cluster compactness and inter-cluster separation. This directly addresses conflicts among agent-generated gradients, common in tasks such as fake news detection or jailbreak classification, and parallels gradient surgery in multi-task learning but for discrete prompts.
Adaptive collaboration. CAAW allocates influence based on agent reliability, reducing noise from weaker gradients while amplifying consistent signals. This is useful when tasks emphasize different reasoning dimensions (e.g., semantic consistency vs. contextual grounding).
Budget-awareness. With lightweight bandit-based selection, MAPGD improves performance-per-token efficiency, achieving higher accuracy with fewer redundant API calls.
Limitations. MAPGD depends on embedding quality and agent reliability; under strong domain shifts, clustering may be unstable and weighting may amplify spurious signals. Future work may strengthen semantic representations and improve agent calibration.
Future directions. Promising avenues include:
  1. (1)
    Cross-task generalization: training reusable agents with transferable clustering and weighting behaviors.
  2. (2)
    Human-in-the-loop fusion: leveraging expert oversight to guide clustering and validate weighting.
  3. (3)
    Hybrid optimization: integrating MAPGD with differentiable prompt tuning to balance interpretability and efficiency.
6. Conclusion
We presented MAPGD: Multi-Agent Prompt Gradient Descent, a novel framework for optimizing prompts in large language models. The key innovations of MAPGD are HCGC, which resolves conflicts through hypersphere-constrained clustering, and CAAW, which adaptively balances the contributions of specialized agents. Together, these mechanisms enable diverse agents to collaborate effectively while maintaining semantic alignment.
Empirically, MAPGD achieves consistent improvements over strong baselines such as ProTeGi, Monte Carlo search, reinforcement learning–based methods, and recent arithmetic reasoning frameworks. On four benchmark classification tasks, MAPGD improves F1 scores by 3–7% on average while reducing token consumption by $\sim$8%, leading to better performance-per-token efficiency. Beyond classification, MAPGD also demonstrates strong transferability to arithmetic reasoning datasets such as GSM8k, AQUARAT, and SVAMP, where it surpasses competitive baselines including PromptWizard. These results confirm that gradient-inspired multi-agent collaboration is broadly applicable across both discriminative and reasoning tasks.
Theoretically, we show that MAPGD achieves a sublinear convergence rate of $O(1/\sqrt{T})$ under standard stochastic approximation assumptions (see Appendix C for details). This result indicates that even in a discrete prompt space, the semantic gradient mechanism—stabilized by HCGC and CAAW—retains the efficiency of classical stochastic gradient descent.
While challenges remain, particularly regarding embedding robustness and LLM feedback reliability, MAPGD provides both a practical toolkit for efficient prompt optimization and a theoretical foundation linking discrete semantic updates with continuous optimization theory. Future work may extend HCGC and CAAW to multimodal prompts, incorporate human preference alignment, and develop cross-domain reusable agents.
In summary, MAPGD demonstrates that multi-agent collaboration, when combined with principled clustering and adaptive weighting, leads to more robust, interpretable, and resource-efficient prompt optimization, with demonstrated effectiveness across classification, reasoning, and beyond.
Appendix A CAAW Parameter Sensitivity
The Channel-Adaptive Agent Weighting module uses a temperature parameter $\lambda$ to control the emphasis on historically effective agents during gradient fusion. To evaluate the robustness of MAPGD with respect to $\lambda$, we conducted experiments on two representative datasets: LIAR(binary) and GSM8K (mathematical reasoning).

### v3

@@ 4.2 Experimental Results
Table 1: Arithmetic reasoning accuracy (%) on GSM8k, AQUARAT, and SVAMP. Best results in bold.
4.3 Ablation Study
To isolate the sources of MAPGD’s gains, we conduct a component-wise ablation by selectively disabling HCGC and CAAW. Table 2 reports four settings (I–IV), where Setting I removes both modules and serves as a minimal baseline. Enabling either module leads to consistent improvements. Introducing HCGC (III vs. I) yields systematic gains, confirming that geometry-aware clustering prior to fusion is essential for reducing semantic interference among pseudo-gradients. CAAW alone (II vs. I) also improves performance, indicating that uniform averaging is suboptimal under varying agent reliability; adaptive reweighting based on validation feedback enhances robustness. The combined setting (IV) consistently outperforms either component alone, suggesting strong complementarity: HCGC resolves directional incompatibility via conflict disentanglement, whereas CAAW addresses reliability mismatch through adaptive credit assignment. Gains are smaller on saturated datasets (e.g., Ethos), consistent with ceiling effects, while conflict-prone tasks benefit more from structured coordination.
Table 2: Ablation study on core MAPGD components. $\checkmark$ indicates the component is enabled. Results are F1 score (mean over 3 runs).
Setting HCGC CAAW Jailbreak Ethos LIAR Sarcasm
I $\times$ $\times$ 0.82 0.91 0.62 0.83
II $\times$ $\checkmark$ 0.84 0.94 0.62 0.87
III $\checkmark$ $\times$ 0.85 0.95 0.63 0.88
IV $\checkmark$ $\checkmark$ 0.88 0.98 0.65 0.91
4.4 Robustness to Dataset Sampling Size
Following the standard ProTeGi evaluation protocol, our main results use 50 development and 150 test examples per dataset to ensure direct comparability. To verify that MAPGD’s gains are not an artifact of small evaluation sets, we additionally evaluate larger subsamples on three representative classification benchmarks.
Table 3 reports results under three sampling regimes (50/150, 100/300, and 200/500). MAPGD consistently outperforms ProTeGi and MC across all datasets and sampling sizes, indicating that its improvements persist as the evaluation set grows.
Dataset Method 50/150 100/300 200/500
Jailbreak ProTeGi 0.82 0.83 0.81
MC 0.76 0.77 0.75
@@ 4.4 Robustness to Dataset Sampling Size
Ethos ProTeGi 0.93 0.94 0.94
MC 0.89 0.90 0.90
[…]
LIAR ProTeGi 0.65 0.62 0.63
MC 0.62 0.63 0.63
[…]
Table 3: Robustness to dataset sampling size (F1 score) under three dev/test splits.
4.5 Token Consumption Analysis
Token efficiency is evaluated under the setup of Section 4.1 by counting both input and output tokens per API call. We report total tokens, number of calls, and performance-per-token. As shown in Table 4, MAPGD attains higher F1 with fewer tokens and calls than ProTeGi, reducing per-iteration token cost by $\sim$8% and improving performance-per-token by >>10%, suggesting reduced redundant queries via HCGC and CAAW.
Metric ProTeGi MAPGD (full)
[…]
Table 4: Token consumption comparison between ProTeGi and MAPGD (averaged over 3 runs). MAPGD achieves lower token usage and higher efficiency.
4.6 Computational Overhead Analysis
We compare runtime overhead against ProTeGi under identical settings. MAPGD shares evaluation, prompt expansion, and bandit selection with ProTeGi, but adds multi-agent gradient generation and embedding-based clustering. As shown in Table 5, the extra overhead mainly comes from multi-agent gradient generation, which scales linearly with the number of agents. Embedding and HCGC incur an almost constant cost, and with four agents the overall overhead remains a small fraction of total runtime.
Component ProTeGi (s) MAPGD (s)
Gradient generation 10.2 37.9
Embedding + HCGC 24.4
Prompt expansion 33.7 28.7
Bandit selection 30.3 26.5
Other 85.2 84.0
[…]
Table 5: Runtime breakdown (in seconds) comparing ProTeGi and MAPGD.
4.7 Agent Number Sensitivity
To motivate the use of four specialized agents, we vary the number of agents $N\in\{2,4,6\}$ under the setup of Section 4.1. For $N=2$, only the Instruction Specialist and Example Curator are used; for $N=6$, two additional style/format agents are added. As shown in Figure 4, average F1 improves substantially from $N=2$ to $N=4$, while $N=6$ yields only marginal gains with higher token cost, indicating diminishing returns. Thus, $N=4$ offers the best effectiveness–efficiency trade-off.
Figure 4: Agent number sensitivity analysis. Increasing agents from 2 to 4 yields clear improvements, while further increasing to 6 brings diminishing returns.
4.8 Experimental Insights
MAPGD consistently outperforms single-agent ProTeGi and other baselines, with the largest gains on LIAR and Jailbreak where refinement signals often conflict. Ablations show HCGC and CAAW are complementary: removing either hurts performance, and removing both approaches the single-agent regime. MAPGD is also more token-efficient than ProTeGi (fewer API calls, >>10% higher performance-per-token) and transfers well to arithmetic reasoning, surpassing strong baselines on GSM8k, AQUARAT, and SVAMP. Appendix H provides a qualitative case study illustrating how HCGC and CAAW resolve conflicting pseudo-gradients.
5 Conclusion
This work introduces MAPGD, a coordinated multi-agent framework for prompt optimization that integrates HCGC and CAAW to resolve conflicting refinement directions and reduce the impact of unreliable agents. Across classification and arithmetic reasoning benchmarks, MAPGD consistently improves over strong baselines while lowering token consumption, and it admits a sublinear convergence guarantee of $O(1/\sqrt{T})$ under standard stochastic approximation assumptions (Appendix G). Overall, MAPGD provides an interpretable, robust, and resource-efficient approach to prompt optimization via structured multi-agent collaboration.
Limitations
MAPGD relies on semantic embeddings to detect conflicts and cluster pseudo-gradients; under domain shift or highly specialized jargon, similarity estimates may become unreliable and degrade coordination quality. The method also depends on reasonably informative agent feedback: if multiple agents are systematically misaligned, adaptive weighting may not fully suppress harmful edits. Finally, compared to single-agent methods, MAPGD introduces extra overhead from multi-agent generation and coordination (embedding/clustering), which may limit deployment under strict latency or budget constraints.
Appendix A Datasets Details
@@
Appendix B Baselines Details
@@
Appendix C Extended Evaluation Budgets Beyond 50 Queries
To verify whether our conclusions depend on the 50-query budget, we further extend the evaluation budget to 60 and 70 queries on three representative datasets (Jailbreak, LIAR, and Sarcasm). We keep the same evaluation protocol and report results in Table 6. Overall, the method ranking remains unchanged, and the additional gains beyond 50 are marginal, suggesting that a 50-query budget is sufficient to capture MAPGD’s performance while preserving fairness with prior work (e.g., ProTeGi).
Dataset Budget ProTeGi MC MAPGD
Jailbreak 50 0.833 0.773 0.880
60 0.836 0.773 0.885
70 0.842 0.773 0.887
LIAR 50 0.660 0.630 0.710
60 0.672 0.630 0.720
70 0.681 0.630 0.725
Sarcasm 50 0.876 0.853 0.913
60 0.879 0.853 0.918
70 0.880 0.853 0.920
Table 6: Performance under extended evaluation budgets (60/70 queries). Gains beyond 50 queries are small and do not change the ranking among methods.
Across these datasets, MAPGD improves by only +0.007 (Jailbreak), +0.015 (LIAR), and +0.007 (Sarcasm) from 50 to 70 queries, confirming a clear diminishing-return trend after approximately 30–40 evaluations, while leaving the main conclusions unchanged.
Appendix D CAAW Parameter Sensitivity
The CAAW module uses a temperature parameter $\lambda$ to control the emphasis on historically effective agents during gradient fusion. To evaluate the robustness of MAPGD with respect to $\lambda$, we conducted experiments on two representative datasets: LIAR(binary) and GSM8K (mathematical reasoning).